Minwoo usually suffers from his dental pain. So, he used to go to two dental clinics ‘Suduki DDS’ and ‘Student Health Dental Care Center’. After getting the ‘Root Canal’ treatments in the clinics many times, Minwoo finds out that the wait time to get the treatment in ‘Suduki DDS’ can be modeled by an exponential random variable with mean time 2 hours and the wait time to get the treatment in ‘Student Health Dental Care Center’ can be modeled by a uniform random variable from 0 to 2 hours. Assume the wait time to get the treatment of these two clinics are independent. Let the wait time to get the treatment in ‘Suduki DDS’ and ‘Student Health Dental Care Center’ be X and Y . Given X < 2Y , find the conditional probability that 2X > Y
Answer #1

Answer:-

Given that:-

Let X & Y be waiting time to use treatement in 'suduki DDS' and student Health Dental care center respectively

Given that X \sim Exp(\lambda) and Y \sim U(0,2)

Then pdf of X and Y are

f_{x}(x) = \lambda e^{-\lambda x} ; x>0X= 1/2& f_{y}(y) = \frac{1}{2} ; 0<y<2

Now since x & y are independent , Then

P(X<2y)=\int^{2}_{0}(\int^{2y}_{0} f_{x}(x) dx)f_{y}(y)dy

=\int^{2}_{0}(\int^{2y}_{0} \lambda e^{-\lambda x} dx)\frac{1}{2}dy

= \frac{1}{2} \int^{1}_{0}|-e^{-\lambda x}|^{2y}_{0} dy

=\frac{1}{2}\int^{2}_{0}(1-e^{-y})dy

=\frac{1}{2}[\int^{2}_{0} dy - \int^{2}_{0}e^{-y}dy]

=\frac{1}{2}[|y|^{2}_{0}+|e^{-y}|^{2}_{0}]

=\frac{1}{2}[2+e^{-2}-1]

=\frac{1+e^{-2}}{2}=0.5677

Hence the conditional probability is

P(2x>y\setminus x<2y)=P(\frac{y}{2}<x<2y,0<y<2)

=\int^{2}_{0}(\int^{2y}_{y/2}f_{x}(x)dx)f_{y}(y)dy

=\frac{1}{2} \int^{2}_{0}(\int^{2y}_{y/2}\lambda e^{-\lambda x}dx)dy

=\frac{1}{2}\int^{2}_{0}|-e^{-\lambda x}|^{2y}_{y/2}dy

=\frac{1}{2}\int^{2}_{0}(e^{-y/4}-e^{-y})dy

= \frac{1}{2}[|-4e^{-y/4}|^{2}_{0}+|e^{-y}|^{2}_{0}]

=\frac{1}{2}[4-4e^{-1/2}+e^{-2}-e^{0}]

= \frac{1}{2}[3+e^{-2}-4e^{-1/2}]

=\frac{1}{2}\times0.709212644\approx 0.3546

Hence P(2x>y\setminus x<2y)=\frac{0.3546}{0.5677}=0.6246